# Permissioned LLMs: Enforcing Access Control in Large Language Models

### v1

@@ 6 Conclusion and Discussion
Limitations. Our approach does not support deep hierarchy of domains with arbitrary overlaps. Another issue we observe is with the scalability beyond a handful of domains. This either leads to severe degradation of utility (as in the case of Activate) or it becomes compute-intensive (for Union). We leave this exploration for future work. We also note some limitations in the experiments that we do not expect to change our key claims. First, we only run one model fine-tuning per parameter setting due to the computation overhead. Second, we do not perform an ablation study on the LoRA rank on fine-tuning. Our preliminary experiments with different ranks suggested limited impact on model utility, so we stick to the default value. For our formalism in § 2, we assume that adversaries do not tamper with their credentials or domain access, otherwise they can gain arbitrary domain information. This is enforced by the enclosing system via authentication.
Related Work. Access control problems in agentic systems can manifest in interesting ways, such as context hijacking [2], and may require further constraining the purview of individual agent contexts. Retrieval Augmented Generation (RAG) systems [22, 31, 50] are also susceptible to the access control problem. However, the access control needs to be enforced in the information retrieval engine of the system [4, 14] and is beyond our work’s scope (although we do provide a formalism for access control in RAG-based systems in Appendix A).
One may draw some parallels between our formalism of response relevance and access advantage metric with prior works on indistinguishability [1, 3, 9, 13] in security and privacy. The mechanisms in this lineage of works are singularly focused on eliminating distinguishability between the effects of different data on computations. In contrast, PermLLM’s objective is to maximize domain separation, which implies maximization of distinguishability – the more pronounced the distinguishability, the more effective is the PermLLM mechanism.
Broader Impacts. We do not foresee any negative societal impact of our work. Our work aims to bolster the security and privacy of individual’s data by enforcing strict access control, such that only people with prior authorization can get access to the information. Our work is applicable to healthcare, finance, and more broadly, enterprise / governance applications that deal with sensitive data of individuals.
@@ 7 Acknowledgements
We would like to thank Pallika Kanani and Dan Roth for fruitful discussions that motivated the problem setup of access control in large language model based applications.
@@ Appendix A Formalizing Access Control for Retrieval Augmented Generation
To empirically quantify response relevance, we can use the same response relevance score, $relv(f_{R}(q),S_{u})$ that quantifies the information gained on data in the security domains $q$’s user $u$ has access to (this is the same set of security domains that mapping $\mathcal{M}$ gives for $u$ for the retriever $R_{C}^{\mathcal{M}}$, i.e. $S_{u}$). Here $R_{C}^{\mathcal{M}}$ retrieves the query context $c\in C$ using mapping $\mathcal{M}$; $c$ is then augmented to the query $q$. We restrict the domain of $relv$ to the real number interval $[0,1]$, where $1$ is the best expected score for relevance. Similar to PermLLM, we define access advantage for PermRAG as follows:
@@ D.2 Data Sets
For our experiments, we require data sets that consist of multiple distinct domains and are possibly not seen by the pretrained models. We use four different data sets, namely, WMDP [23], GPQA [32], SimpleQA [40], and RCV1 [21]. While the first three data sets were collected after the pretraining cutoff dates for Llama-3.1-8B and Mistral-0.1-7B, RCV1 is an older data set and hence we do not know if it was used in pretraining. However, we observe a high initial training loss on this data set, thereby indicating that it was either not used in pretraining or was catastrophically forgotten by the models, allowing for a gradual reduction in training loss during our fine-tuning (see Figure 7). Table 6 shows the data set details, along with the generalization gap (test loss - train loss) for different approaches of fine-tuning the models on these data sets. See Figure 4, Figure 5, Figure 6, and Figure 7 for complete training and test loss trajectories across different data sets.
@@ D.3 Model Utility Evaluation
We use four metrics to evaluate the utility of the model generations: Bleurt Score ($bluert$), Bert F1-Score ($bert$), Sacrebleu Score ($bleu$) and Verbatim Accuracy ($acc$). These metrics measure how similar the generated text is to the ground truth. $bleurt$ and $bert$ measure the semantic similarity, $bleu$ measures the fraction of common n-grams, and $acc$ gives a binary decision of whether the generated text verbatim matches the ground truth. All the metrics lie in a [0,1] range, where values close to 1 indicate high model utility.
We check the utility of Activate to determine if tuning different LoRA adapters for each security domain leads to acceptable model utility. To that end, we show in Table 7 the utility of Llama-3.1-8B models fine-tuned on different data sets with the three approaches: full fine-tuning, LoRA fine-tuning and our PermLLM. We do not report the $bleu$ score for WMDP as it is a multi-choice question-answering task where model only has to generate a single token. $bleu$ requires generating at least four tokens. Our approach achieves similar or better utility on the training set compared to the LoRA approach. On the test set, our approach achieves similar utility to LoRA for most of the data sets, except for SimpleQA where LoRA performs better. This is because SimpleQA has more domains (10 in total), thus each of our individual domain adapter sees only a fraction of data of what LoRA approach’s adapter sees (given that SimpleQA is already a small data set). We expect the performance of our domain-specific adapters to increase as the data set size increases. Full fine-tuning is highly sensitive to training hyper-parameters, and as a result it either completely overfits on training set to achieve high utility (e.g., on SimpleQA and RCV1), or it underfits and achieves low utility (e.g., on WMDP and GPQA). We observe similar results for Mistral-0.1-7B models (see Table 8).
@@
WMDP $bleurt$ $0.74\pm 0.06$ $0.74\pm 0.06$ $0.90\pm 0.08$ $0.85\pm 0.08$ $0.92\pm 0.08$ $0.82\pm 0.06$
$bert$ $0.89\pm 0.03$ $0.89\pm 0.03$ $0.96\pm 0.03$ $0.94\pm 0.03$ $0.97\pm 0.03$ $0.93\pm 0.03$
$acc$ $0.26\pm 0.07$ $0.27\pm 0.07$ $0.76\pm 0.20$ $0.60\pm 0.20$ $0.84\pm 0.22$ $0.49\pm 0.15$
GPQA $bleu$ $0.26\pm 0.02$ $0.05\pm 0.03$ $0.45\pm 0.12$ $0.10\pm 0.05$ $0.39\pm 0.20$ $0.10\pm 0.04$
$bleurt$ $0.53\pm 0.05$ $0.39\pm 0.05$ $0.64\pm 0.09$ $0.46\pm 0.07$ $0.62\pm 0.11$ $0.47\pm 0.07$
$bert$ $0.67\pm 0.06$ $0.59\pm 0.05$ $0.77\pm 0.08$ $0.67\pm 0.05$ $0.75\pm 0.09$ $0.67\pm 0.05$
$acc$ $0.24\pm 0.06$ $0.02\pm 0.03$ $0.32\pm 0.05$ $0.05\pm 0.05$ $0.31\pm 0.09$ $0.04\pm 0.05$
SimpleQA $bleu$ $0.80\pm 0.06$ $0.34\pm 0.11$ $0.65\pm 0.06$ $0.29\pm 0.08$ $0.67\pm 0.10$ $0.09\pm 0.04$
$bleurt$ $0.86\pm 0.03$ $0.58\pm 0.05$ $0.80\pm 0.02$ $0.61\pm 0.02$ $0.82\pm 0.04$ $0.53\pm 0.04$
$bert$ $0.96\pm 0.01$ $0.84\pm 0.02$ $0.94\pm 0.01$ $0.86\pm 0.01$ $0.95\pm 0.02$ $0.82\pm 0.03$
$acc$ $0.68\pm 0.10$ $0.20\pm 0.12$ $0.52\pm 0.07$ $0.17\pm 0.07$ $0.55\pm 0.13$ $0.02\pm 0.02$
RCV1 $bleu$ $0.75\pm 0.08$ $0.14\pm 0.08$ $0.22\pm 0.10$ $0.16\pm 0.08$ $0.27\pm 0.10$ $0.16\pm 0.08$
$bleurt$ $0.88\pm 0.04$ $0.46\pm 0.12$ $0.57\pm 0.13$ $0.49\pm 0.11$ $0.62\pm 0.13$ $0.50\pm 0.12$
$bert$ $0.94\pm 0.03$ $0.67\pm 0.09$ $0.75\pm 0.08$ $0.70\pm 0.07$ $0.78\pm 0.08$ $0.70\pm 0.08$
$acc$ $0.78\pm 0.06$ $0.16\pm 0.10$ $0.27\pm 0.14$ $0.17\pm 0.10$ $0.31\pm 0.15$ $0.18\pm 0.10$
@@
WMDP Bleurt $0.95\pm 0.01$ $0.82\pm 0.03$ $0.96\pm 0.02$ $0.87\pm 0.03$ $0.96\pm 0.01$ $0.86\pm 0.03$
Bert $0.98\pm 0.01$ $0.92\pm 0.02$ $0.99\pm 0.01$ $0.94\pm 0.02$ $0.99\pm 0.01$ $0.94\pm 0.02$
Acc $0.88\pm 0.04$ $0.46\pm 0.14$ $0.92\pm 0.07$ $0.60\pm 0.09$ $0.93\pm 0.04$ $0.58\pm 0.11$
GPQA Bleu $0.46\pm 0.03$ $0.06\pm 0.05$ $0.35\pm 0.08$ $0.11\pm 0.07$ $0.55\pm 0.18$ $0.13\pm 0.06$
Bleurt $0.65\pm 0.04$ $0.42\pm 0.08$ $0.59\pm 0.09$ $0.47\pm 0.06$ $0.67\pm 0.09$ $0.47\pm 0.08$
Bert $0.75\pm 0.05$ $0.62\pm 0.07$ $0.73\pm 0.08$ $0.68\pm 0.05$ $0.79\pm 0.08$ $0.66\pm 0.09$
Acc $0.38\pm 0.04$ $0.04\pm 0.05$ $0.24\pm 0.04$ $0.05\pm 0.06$ $0.40\pm 0.09$ $0.08\pm 0.02$
SimpleQA Bleu $0.94\pm 0.02$ $0.36\pm 0.11$ $0.73\pm 0.06$ $0.34\pm 0.09$ $0.70\pm 0.13$ $0.10\pm 0.04$
Bleurt $0.94\pm 0.01$ $0.60\pm 0.04$ $0.84\pm 0.03$ $0.62\pm 0.03$ $0.83\pm 0.06$ $0.52\pm 0.04$
Bert $0.99\pm 0.01$ $0.85\pm 0.02$ $0.96\pm 0.01$ $0.87\pm 0.01$ $0.95\pm 0.03$ $0.82\pm 0.03$
Acc $0.91\pm 0.04$ $0.23\pm 0.12$ $0.62\pm 0.08$ $0.20\pm 0.10$ $0.60\pm 0.16$ $0.03\pm 0.02$
RCV1 Bleu $0.92\pm 0.06$ $0.17\pm 0.09$ $0.28\pm 0.13$ $0.20\pm 0.10$ $0.37\pm 0.14$ $0.19\pm 0.09$
Bleurt $0.93\pm 0.02$ $0.48\pm 0.12$ $0.60\pm 0.13$ $0.51\pm 0.12$ $0.66\pm 0.12$ $0.50\pm 0.12$
Bert $0.98\pm 0.02$ $0.69\pm 0.08$ $0.78\pm 0.09$ $0.71\pm 0.08$ $0.81\pm 0.08$ $0.71\pm 0.08$
Acc $0.92\pm 0.03$ $0.19\pm 0.11$ $0.31\pm 0.15$ $0.20\pm 0.11$ $0.38\pm 0.17$ $0.19\pm 0.10$
@@
Appendix E MIAs against LLMs

### v2

@@ 6 Conclusion and Discussion
Limitations. Our approach does not support deep hierarchy of domains with arbitrary overlaps. Another issue we observe is with the scalability beyond a handful of domains. This either leads to severe degradation of utility (as in the case of Activate) or it becomes compute-intensive (for Union). We leave this exploration for future work. We also note some limitations in the experiments that we do not expect to change our key claims. First, we only run one model fine-tuning per parameter setting due to the computation overhead. Second, we use the default value for LoRA rank as our preliminary experiments with different ranks suggested limited impact on model utility. For our formalism in § 2, we assume that adversaries do not tamper with their credentials or domain access, otherwise they can gain arbitrary domain information. This is enforced by the enclosing system via authentication. Finally, we note that our DDI metrics are only as good as the best possible MIA.
Related Work. Access control problems in agentic systems can manifest in interesting ways, such as context hijacking [2], and may require further constraining the purview of individual agent contexts. Retrieval Augmented Generation (RAG) systems [23, 32, 51] are also susceptible to the access control problem. However, the access control needs to be enforced in the information retrieval engine of the system [4, 14] and is beyond our work’s scope (although we do provide a formalism for access control in RAG-based systems in Appendix A). One may draw some parallels between our formalism of response relevance and access advantage metric with prior works on indistinguishability [1, 3, 9, 13] in security and privacy. The mechanisms in this lineage of works are singularly focused on eliminating distinguishability between the effects of different data on computations. In contrast, PermLLM’s objective is to maximize domain separation, which implies maximization of distinguishability.
Broader Impacts. Our work aims to bolster the security and privacy of individual’s data by enforcing strict access control, such that only people with prior authorization can get access to the information. Our work is applicable to healthcare, finance, and more broadly, enterprise / governance applications that deal with sensitive data of individuals.
@@ 7 Acknowledgements
We would like to thank Pallika Kanani and Dan Roth for fruitful discussions that motivated the problem setup of access control in large language model based applications. We would also like to thank David Evans and the anonymous reviewers for providing helpful feedback on the paper.
@@ Appendix A Formalizing Access Control for Retrieval Augmented Generation
To empirically quantify response relevance, we can use the same response relevance score, $relv(f_{R}(q),S_{u})$ that quantifies the information gained on data in the security domains $q$’s user $u$ has access to (this is the same set of security domains that mapping $\mathcal{M}$ gives for $u$ for the retriever $R_{C}^{\mathcal{M}}$, i.e. $S_{u}$). Here $R_{C}^{\mathcal{M}}$ retrieves the query context $c\in C$ using mapping $\mathcal{M}$; $c$ is then augmented to the query $q$. We restrict the domain of relvrelv to the real number interval $[0,1]$, where $1$ is the best expected score for relevance. Similar to PermLLM, we define access advantage for PermRAG as follows:
@@ D.2 Data Sets
For our experiments, we require data sets that consist of multiple distinct domains and are possibly not seen by the pretrained models. We use five different data sets, namely, WMDP [24], GPQA [33], SimpleQA [41], RCV1 [22], and PubMedQA [20]. While the first three data sets were collected after the pretraining cutoff dates for Llama-3.1-8B and Mistral-0.1-7B, RCV1 is an older data set and hence we do not know if it was used in pretraining. However, we observe a high initial training loss on this data set, thereby indicating that it was either not used in pretraining or was catastrophically forgotten by the models, allowing for a gradual reduction in training loss during our fine-tuning (see Figure 7). Table 6 shows the data set details, along with the generalization gap (test loss - train loss) for different approaches of fine-tuning the models on these data sets. See Figure 4, Figure 5, Figure 6, Figure 7, and Figure 8 for complete training and test loss trajectories across different data sets.
@@
Figure 8: Comparing model loss on PubMedQA data set.
@@ D.3 Model Utility Evaluation
We use four metrics to evaluate the utility of the model generations: Bleurt Score (bluertbluert), Bert F1-Score (bertbert), Sacrebleu Score (bleubleu) and Verbatim Accuracy (accacc). These metrics measure how similar the generated text is to the ground truth. bleurtbleurt and bertbert measure the semantic similarity, bleubleu measures the fraction of common n-grams, and accacc gives a binary decision of whether the generated text verbatim matches the ground truth. All the metrics lie in a [0,1] range, where values close to 1 indicate high model utility.
We check the utility of Activate to determine if tuning different LoRA adapters for each security domain leads to acceptable model utility. To that end, we show in Table 7 the utility of Llama-3.1-8B models fine-tuned on different data sets with the three approaches: full fine-tuning, LoRA fine-tuning and our PermLLM. We do not report the bleubleu score for WMDP as it is a multi-choice question-answering task where model only has to generate a single token. bleubleu requires generating at least four tokens. Our approach achieves similar or better utility on the training set compared to the LoRA approach. On the test set, our approach achieves similar utility to LoRA for most of the data sets, except for SimpleQA where LoRA performs better. This is because SimpleQA has more domains (10 in total), thus each of our individual domain adapter sees only a fraction of data of what LoRA approach’s adapter sees (given that SimpleQA is already a small data set). We expect the performance of our domain-specific adapters to increase as the data set size increases. Full fine-tuning is highly sensitive to training hyper-parameters, and as a result it either completely overfits on training set to achieve high utility (e.g., on SimpleQA and RCV1), or it underfits and achieves low utility (e.g., on WMDP and GPQA). We observe similar results for Mistral-0.1-7B models (see Table 8).
@@
WMDP bleurtbleurt $0.74\pm 0.06$ $0.74\pm 0.06$ $0.90\pm 0.08$ $0.85\pm 0.08$ $0.92\pm 0.08$ $0.82\pm 0.06$
bertbert $0.89\pm 0.03$ $0.89\pm 0.03$ $0.96\pm 0.03$ $0.94\pm 0.03$ $0.97\pm 0.03$ $0.93\pm 0.03$
accacc $0.26\pm 0.07$ $0.27\pm 0.07$ $0.76\pm 0.20$ $0.60\pm 0.20$ $0.84\pm 0.22$ $0.49\pm 0.15$
GPQA bleubleu $26\pm 0.02 0.0$$05\pm 0.03$ $0.45\pm 0.12$ $0.10\pm 0.05$ $0.39\pm 0.20$ $0.10\pm 0.04$
bleurtbleurt $0.53\pm 0.05$ $0.39\pm 0.05$ $0.64\pm 0.09$ $0.46\pm 0.07$ $0.62\pm 0.11$ $0.47\pm 0.07$
bertbert $0.67\pm 0.06$ $0.59\pm 0.05$ $0.77\pm 0.08$ $0.67\pm 0.05$ $0.75\pm 0.09$ $0.67\pm 0.05$
accacc $24\pm 0.06 0.0$$02\pm 0.03$ $32\pm 0.05 0.0$$05\pm 0.05$ $31\pm 0.09 0.0$$04\pm 0.05$
SimpleQA bleubleu $0.80\pm 0.06$ $0.34\pm 0.11$ $0.65\pm 0.06$ $0.29\pm 0.08$ $67\pm 0.10 0.0$$09\pm 0.04$
bleurtbleurt $0.86\pm 0.03$ $0.58\pm 0.05$ $0.80\pm 0.02$ $0.61\pm 0.02$ $0.82\pm 0.04$ $0.53\pm 0.04$
bertbert $0.96\pm 0.01$ $0.84\pm 0.02$ $0.94\pm 0.01$ $0.86\pm 0.01$ $0.95\pm 0.02$ $0.82\pm 0.03$
accacc $0.68\pm 0.10$ $0.20\pm 0.12$ $0.52\pm 0.07$ $0.17\pm 0.07$ $55\pm 0.13 0.0$$02\pm 0.02$
RCV1 bleubleu $0.75\pm 0.08$ $0.14\pm 0.08$ $0.22\pm 0.10$ $0.16\pm 0.08$ $0.27\pm 0.10$ $0.16\pm 0.08$
bleurtbleurt $0.88\pm 0.04$ $0.46\pm 0.12$ $0.57\pm 0.13$ $0.49\pm 0.11$ $0.62\pm 0.13$ $0.50\pm 0.12$
bertbert $0.94\pm 0.03$ $0.67\pm 0.09$ $0.75\pm 0.08$ $0.70\pm 0.07$ $0.78\pm 0.08$ $0.70\pm 0.08$
accacc $0.78\pm 0.06$ $0.16\pm 0.10$ $0.27\pm 0.14$ $0.17\pm 0.10$ $0.31\pm 0.15$ $0.18\pm 0.10$
PubMedQA bleubleu $71\pm 0.05 0.0$$7\pm 0.01 0.0$$9\pm 0.01 0.0$$09\pm 0.01$ $10\pm 0.02 0.0$$09\pm 0.01$
bleurtbleurt $0.77\pm 0.03$ $0.38\pm 0.01$ $0.40\pm 0.01$ $0.40\pm 0.01$ $0.42\pm 0.01$ $0.40\pm 0.02$
bertbert $0.90\pm 0.02$ $0.64\pm 0.02$ $0.68\pm 0.01$ $0.68\pm 0.01$ $0.69\pm 0.02$ $0.67\pm 0.02$
accacc - - - - - -
@@
WMDP bleurtbleurt $0.95\pm 0.01$ $0.82\pm 0.03$ $0.96\pm 0.02$ $0.87\pm 0.03$ $0.96\pm 0.01$ $0.86\pm 0.03$
bertbert $0.98\pm 0.01$ $2\pm 0.02 0.99±0$$9\pm 0.01$ $4\pm 0.02 0.99±0$$9\pm 0.01$ $0.94\pm 0.02$
accacc $0.88\pm 0.04$ $0.46\pm 0.14$ $0.92\pm 0.07$ $0.60\pm 0.09$ $0.93\pm 0.04$ $0.58\pm 0.11$
GPQA bleubleu $46\pm 0.03 0.0$$06\pm 0.05$ $0.35\pm 0.08$ $0.11\pm 0.07$ $0.55\pm 0.18$ $0.13\pm 0.06$
bleurtbleurt $0.65\pm 0.04$ $0.42\pm 0.08$ $0.59\pm 0.09$ $0.47\pm 0.06$ $0.67\pm 0.09$ $0.47\pm 0.08$
bertbert $0.75\pm 0.05$ $0.62\pm 0.07$ $0.73\pm 0.08$ $0.68\pm 0.05$ $0.79\pm 0.08$ $0.66\pm 0.09$
accacc $38\pm 0.04 0.0$$04\pm 0.05$ $24\pm 0.04 0.0$$05\pm 0.06$ $40\pm 0.09 0.0$$08\pm 0.02$
SimpleQA bleubleu $0.94\pm 0.02$ $0.36\pm 0.11$ $0.73\pm 0.06$ $0.34\pm 0.09$ $0.70\pm 0.13$ $0.10\pm 0.04$
bleurtbleurt $0.94\pm 0.01$ $0.60\pm 0.04$ $0.84\pm 0.03$ $0.62\pm 0.03$ $0.83\pm 0.06$ $0.52\pm 0.04$
bertbert $0.99\pm 0.01$ $0.85\pm 0.02$ $0.96\pm 0.01$ $0.87\pm 0.01$ $0.95\pm 0.03$ $0.82\pm 0.03$
accacc $0.91\pm 0.04$ $0.23\pm 0.12$ $0.62\pm 0.08$ $0.20\pm 0.10$ $60\pm 0.16 0.0$$03\pm 0.02$
RCV1 bleubleu $0.92\pm 0.06$ $0.17\pm 0.09$ $0.28\pm 0.13$ $0.20\pm 0.10$ $0.37\pm 0.14$ $0.19\pm 0.09$
bleurtbleurt $0.93\pm 0.02$ $0.48\pm 0.12$ $0.60\pm 0.13$ $0.51\pm 0.12$ $0.66\pm 0.12$ $0.50\pm 0.12$
bertbert $0.98\pm 0.02$ $0.69\pm 0.08$ $0.78\pm 0.09$ $0.71\pm 0.08$ $0.81\pm 0.08$ $0.71\pm 0.08$
cccc $0.92\pm 0.03$ $0.19\pm 0.11$ $0.31\pm 0.15$ $0.20\pm 0.11$ $0.38\pm 0.17$ $0.19\pm 0.10$
PubMedQA bleubleu $75\pm 0.04 0.0$$8\pm 0.01 0.0$$9\pm 0.01 0.0$$08\pm 0.01$ $11\pm 0.02 0.0$$08\pm 0.01$
bleurtbleurt $0.80\pm 0.03$ $0.39\pm 0.01$ $0.41\pm 0.01$ $0.41\pm 0.01$ $0.43\pm 0.02$ $0.41\pm 0.01$
bertbert $0.92\pm 0.01$ $0.65\pm 0.02$ $0.69\pm 0.01$ $0.68\pm 0.02$ $0.70\pm 0.02$ $0.68\pm 0.01$
accacc - - - - - -
@@
Appendix E Prompt-Based Access Control
Recent works [8, 25] have proposed enforcing some form of access control in system prompts, however we note that they do not provide absolute access control and are vulnerable to jailbreaking prompts. Regardless, we implement prompt-based access control as a baseline where each query is tagged with a prompt prefix (e.g., “use domain 1”) and the rest of the fine-tuning pipeline is similar to LoRA fine-tuning. We add the relevant prompt prefixes during both model fine-tuning and inference. The models fine-tuned with prompt-based access control achieve similar training and test loss to that of LoRA fine-tuning across all the data sets, as shown in Figure 4, Figure 5, Figure 6, Figure 7, and Figure 8. However, this baseline fails to provide any meaningful access control, even when a user has access to only one security domain as shown in Figure 10 and Table 9. As shown in the figure and table, the utility gap index is close to zero and DDI scores are close to random guessing across all the data sets for both Llama and Mistral models fine-tuned with prompt-based access control. The reason is that the prompt prefix for different domains only differ in one or two tokens and hence the model tends to ignore this difference and continues generating responses even for domains the user has no access to. Exploring different prompt structures might lead to better access control but is beyond the scope of this work. We observe a similar trend when the user has access to multiple security domains as shown in Figure 11 for SimpleQA data set.
Figure 10: Utility Gap Index, $\Delta_{U}$ ($mean\pm std$) for prompt-based access control baseline when user has access to one security domain.
Figure 11: Utility Gap Index, $\Delta_{U}$ ($mean\pm std$) for prompt-based access control baseline on different models fine-tuned on SimpleQA when user has access to multiple security domains.
Table 9: DDI values for prompt-based access control baseline when user has access to one security domain.
MIA Llama-3.1-8B Mistral-0.1-7B
auc-roc tpr@1%fpr tpr@5%fpr auc-roc tpr@1%fpr tpr@5%fpr
WMDP Loss $53\pm 0.02 0.0$$2\pm 0.01 0.0$$06\pm 0.01$ $54\pm 0.02 0.0$$2\pm 0.01 0.0$$07\pm 0.02$
ZLIB $52\pm 0.01 0.0$$1\pm 0.01 0.0$$06\pm 0.01$ $52\pm 0.01 0.0$$1\pm 0.01 0.0$$06\pm 0.01$
Mink $53\pm 0.03 0.0$$2\pm 0.02 0.0$$08\pm 0.03$ $53\pm 0.01 0.0$$2\pm 0.01 0.0$$06\pm 0.01$
Mink++ $55\pm 0.06 0.0$$2\pm 0.03 0.0$$08\pm 0.05$ $52\pm 0.03 0.0$$1\pm 0.00 0.0$$05\pm 0.01$
Ref $53\pm 0.02 0.0$$2\pm 0.01 0.0$$06\pm 0.00$ $53\pm 0.01 0.0$$1\pm 0.01 0.0$$06\pm 0.01$
GPQA Loss $55\pm 0.02 0.0$$2\pm 0.00 0.0$$06\pm 0.10$ $56\pm 0.03 0.0$$03\pm 0.01$ $0.12\pm 0.04$
ZLIB $54\pm 0.02 0.0$$2\pm 0.00 0.0$$07\pm 0.01$ $54\pm 0.02 0.0$$3\pm 0.01 0.0$$08\pm 0.02$
Mink $57\pm 0.05 0.0$$02\pm 0.01$ $0.12\pm 0.02$ $59\pm 0.07 0.0$$05\pm 0.06$ $0.13\pm 0.07$
Mink++ $54\pm 0.10 0.0$$05\pm 0.06$ $0.12\pm 0.08$ $55\pm 0.12 0.0$$06\pm 0.06$ $0.12\pm 0.09$
Ref $57\pm 0.02 0.0$$04\pm 0.02$ $0.13\pm 0.05$ $56\pm 0.03 0.0$$03\pm 0.02$ $0.13\pm 0.07$
SimpleQA Loss $53\pm 0.25 0.0$$08\pm 0.14$ $0.16\pm 0.20$ $55\pm 0.22 0.0$$09\pm 0.15$ $0.16\pm 0.20$
ZLIB $52\pm 0.16 0.0$$4\pm 0.05 0.0$$09\pm 0.09$ $53\pm 0.14 0.0$$3\pm 0.03 0.0$$09\pm 0.08$
Mink $52\pm 0.28 0.0$$09\pm 0.15$ $0.17\pm 0.22$ $55\pm 0.22 0.0$$09\pm 0.15$ $0.17\pm 0.20$
Mink++ $0.50\pm 0.43$ $0.31\pm 0.40$ $0.36\pm 0.43$ $0.52\pm 0.35$ $0.22\pm 0.33$ $0.28\pm 0.35$
Ref $53\pm 0.22 0.0$$03\pm 0.03$ $0.11\pm 0.12$ $54\pm 0.15 0.0$$4\pm 0.06 0.0$$09\pm 0.09$
RCV1 Loss $50\pm 0.02 0.0$$1\pm 0.00 0.0$$05\pm 0.01$ $50\pm 0.01 0.0$$1\pm 0.00 0.0$$05\pm 0.00$
ZLIB $50\pm 0.01 0.0$$1\pm 0.00 0.0$$05\pm 0.02$ $50\pm 0.00 0.0$$1\pm 0.00 0.0$$05\pm 0.00$
Mink $50\pm 0.04 0.0$$1\pm 0.00 0.0$$05\pm 0.01$ $50\pm 0.01 0.0$$1\pm 0.02 0.0$$05\pm 0.01$
Mink++ $50\pm 0.05 0.0$$1\pm 0.01 0.0$$05\pm 0.01$ $50\pm 0.04 0.0$$1\pm 0.00 0.0$$05\pm 0.01$
Ref $50\pm 0.01 0.0$$1\pm 0.01 0.0$$05\pm 0.01$ $50\pm 0.01 0.0$$1\pm 0.00 00$$0.5\pm 0.00$
PubMedQA Loss $50\pm 0.00 0.0$$1\pm 0.00 0.0$$05\pm 0.00$ $50\pm 0.00 0.0$$1\pm 0.00 0.0$$05\pm 0.00$
ZLIB $50\pm 0.00 0.0$$1\pm 0.00 0.0$$05\pm 0.00$ $50\pm 0.00 0.0$$1\pm 0.00 0.0$$05\pm 0.00$
Mink $50\pm 0.01 0.0$$1\pm 0.00 0.0$$05\pm 0.00$ $50\pm 0.01 0.0$$1\pm 0.00 0.0$$05\pm 0.01$
Mink++ $50\pm 0.01 0.0$$1\pm 0.00 0.0$$05\pm 0.00$ $50\pm 0.01 0.0$$1\pm 0.00 0.0$$05\pm 0.00$
Ref $50\pm 0.03 0.0$$1\pm 0.00 0.0$$05\pm 0.01$ $50\pm 0.03 0.0$$1\pm 0.00 0.0$$05\pm 0.01$
Appendix F MIAs against LLMs